# Lrq-Fact: LLM-Generated Relevant Questions for Multimodal Fact-Checking

### v1

@@ A Appendix
Language Model with 7B Parameter
InstructBLIP [11] Vicuna-7B 14.7 30.8 13.2 8.1 12.8 33.1 8.2 8.9
VILA [39] Llama2-7B 41.2 35.0 50.0 70.0 26.0 38.2 52.1 67.4
BLIP-2 [34] FlanT5-XL 41.2 35.0 50.0 70.0 25.9 38.3 53.2 68.1
LLaVA-1.6 [42] Vicuna-7B 48.1 48.2 48.5 59.5 37.2 41.2 47.5 54.3
Language Model with 13B Parameter
VILA [39] Llama2-13B 41.1 35.0 50.0 70.0 43.1 32.0 49.2 67.4
InstructBLIP [11] Vicuna-13B 41.1 35.0 49.9 69.9 34.3 48.2 45.7 64.9
BLIP-2 [34] FlanT5-XXL 31.6 63.4 53.6 35.5 30.2 47.2 33.3 31.8
LLaVA-1.6 [42] Vicuna-13B 41.1 35.0 50.0 69.7 34.1 31.2 52.6 65.2
Language Model with 34B Parameter
LLaVA-1.6 [42] Nous-Hermes-2-Yi-34B 62.9 67.1 70.0 63.4 57.1 63.9 67.2 59.3
ChatGPT and Llama3.1
GPT-4V-1.7T [53] ChatGPT 72.3 72.1 72.8 75.6 72.1 74.6 71.8 72.1
Paligemma-3B [6] Llama3.1-70B 72.8 72.1 74.1 75.4 70.9 71.2 72.8 73.1
Paligemma-3B [6] ChatGPT 74.3 76.2 73.1 77.9 72.1 74.3 72.3 74.8
The detailed results are presented in Table 4. Our findings highlight that current models, including both open-source models and GPT-4V, face significant challenges in binary classification tasks on these datasets. Even the advanced GPT-4V model achieves only a 72.3% F1 score with the standard prompt. In contrast, our proposed framework, Lrq-Fact, demonstrates substantial improvements. For example, the BLIP2-FlanT5-XXL model shows a remarkable 1% increase in F1 score when integrated into our framework. This improvement can be attributed to the effective incorporation of rationales, actions, and tool usage, which enhances multimodal understanding in the context of misinformation detection.
A.2 Details of Existing Single-Source Detectors
In this section, we provide detailed descriptions of the single-source detectors used as baselines in our experiments. These detectors represent state-of-the-art models for detecting misinformation in specific modalities, including textual and visual content.
FakeingFakeNews.
FakeingFakeNews [21] is designed to detect textual fake news, focusing particularly on human-written misinformation. It introduces an innovative method for generating training data, leveraging well-known techniques used in human-authored propaganda. The framework uses the ROBERTA [46] model as its backbone and is trained on the PROPANEWS dataset. In our experiments, we applied the ROBERTA detector with its default settings, as provided in the FakeingFakeNews framework.
CNNSpot.
CNNSpot [76] specializes in detecting artificial images generated by models. It utilizes the ResNet-50 model as its classifier backbone. CNNSpot highlights the benefits of data augmentation, such as JPEG compression and Gaussian blur, which improve the detector’s generalization capabilities. In our study, we used the pre-trained CNNSpot model with default hyperparameters to detect visual veracity distortions.
UnivFD.
UnivFD [52] is a versatile fake image detector that operates in a feature space not explicitly trained to differentiate between real and fake images. By accessing the feature space of a pre-trained vision-language model, UnivFD applies a nearest-neighbor approach to detect fake images from various sources. The use of a large pre-trained model enables smoother decision boundaries, enhancing generalization. We utilized the pre-trained UnivFD model with default settings to perform visual veracity distortion detection in our experiments.
LNP.
LNP [40] employs a denoising model to extract noise patterns from spatial images, then analyzes the frequency domain of these noise patterns to detect fake images. It uses the ResNet-50 model as its backbone for classification. In our study, we used the pre-trained LNP detector with default hyperparameters to carry out visual veracity distortion detection.
HAMMER.
HAMMER [61] is a multimodal detector developed to identify multimedia manipulation. Built on the pre-trained vision-language model ALBEF [35], it uses two unimodal encoders and a multimodal aggregator. HAMMER employs hierarchical manipulation reasoning: shallow reasoning checks semantic alignment between image and text embeddings, while deep reasoning performs cross-modal fusion for manipulation detection. In this work, we used the off-the-shelf HAMMER detector with default hyperparameters to detect cross-modal consistency distortion.
FakeNewsGPT4.
FakeNewsGPT4 [44] is based on large vision-language models (LVLMs) and is designed to detect multimodal fake news. It integrates two types of forgery-specific knowledge: semantic correlation and artifact tracking. The model enhances LVLMs by incorporating these two elements, utilizing a multi-level cross-modal reasoning module for semantic correlation and a dual-branch fine-grained verification model for unimodal localized detail comprehension. In our experiments, we used the pre-trained FakeNewsGPT4 with default hyperparameters to detect cross-modal consistency distortion.
A.3 Example Analysis
@@
The questions presented in this case are well-formed and reliable. They are designed to extract key details from both the image and the text, ensuring comprehensive verification. The image-focused questions effectively ask about the setting and identity, which helps in confirming whether the person and location in the image match the article’s claims. The text-based questions aim to validate the timeline and factual details, ensuring a consistent narrative. These questions are precise and structured to get the best possible answers, making them a solid mechanism for cross-verifying facts.
@@
The questions in this case are reliable and appropriately structured to identify discrepancies between the image and the text. The image-focused questions ask about the architectural style and contextual clues from the image, while the text-based questions explore the factual accuracy of the claim that this lighthouse is haunted and located in Greece. The questions provide a good framework for fact-checking by encouraging thorough scrutiny of both visual and textual elements. This ensures that any distortions or misrepresentations in the article are effectively highlighted, making the questions a valuable tool for getting to the truth.
@@
The questions are well-crafted to reveal any visual inconsistencies. The image-focused questions ask about the color and reality of the clock tower, which are key to identifying that the clock tower has been digitally altered. The text-based questions, which probe the existence of such a clock tower in real life, also help uncover discrepancies. These questions are reliable and precise, aimed at extracting the best possible answers and guiding the evaluation of the article’s claims against the evidence provided by the image.
@@
The questions presented are well-designed to highlight the inconsistency between the text and the image. The image-focused questions ask about the scene and the object the girl is holding, providing clear answers that reveal the mismatch. The text-based questions further confirm this by addressing the article’s lack of accurate description. These questions are well-structured and reliable, allowing for an in-depth examination of both the image and the text to expose cross-modal discrepancies. They guide the analysis toward the best possible answers by focusing on the key elements that need verification.
A.4 Instruct Prompt for Lrq-Fact
@@
Image-Focused Questions Prompt. This stage generates relevant image-focused questions designed to verify the accuracy, authenticity, and relevance of the visual content in relation to the article. These questions help clarify the image content and assess its relation to the text. The specific prompt for generating these questions is illustrated in Figure 9.
Image-Focused Answers Prompt. After generating the image-focused questions, this prompt helps in generating answers that analyze the visual content directly from the image. These answers are based on the key elements and actions identified in the image, ensuring that the responses are relevant and insightful. The specific prompt for this is shown in 10.
Text-Focused Questions Prompt. To critically assess the factual claims in the text, this prompt generates relevant questions targeting specific elements such as dates, names, locations, and events. The generated questions aim to challenge the accuracy of the claims made in the article. The specific prompt used for text-focused questions is shown in Figure 11.
Text-Focused Answers Prompt. After generating the text-focused questions, this prompt helps generate accurate answers using the model’s built-in knowledge. These answers verify factual accuracy and challenge any unsupported claims in the article. The specific prompt for this is illustrated in Figure 12.

### v2

@@ 5.7 Experiment Analysis
w/ Visual & Textual FCQs (w/o RAG)
62.8+22.6% 61+8.1% 47.7+15.7% 54.8+2.2% - -
[…]
w/ Visual & Textual FCQs (w/ RAG)
64.2+25.3% 64.8+14.8% 48.6+17.9% 62.1+15.8% 73.2+10.5% 71.5+9.4%
Lrq-Fact (VLM & LLM: GPT-4o) 49.2 60.9 39.9 55.9 72.5 71.2
     w/ Textual FCQs 64.3+30.7% 62.1+2.0% 40.4+1.3% 54.3-2.9% 71.1-1.9% 68.7-3.5%
[…]
w/ Visual & Textual FCQs (w/o RAG)
66.5+35.2% 65.5+7.5% 45.8+14.8% 58.0+3.8% - -
[…]
w/ Visual & Textual FCQs (w/ RAG)
71.6+45.5% 70.8+16.3% 49.2+23.3% 62.3+11.5% 75.2+3.7% 73.1+2.7%
Table 3: Ablation study result. The subscript values indicate the percentage improvement over the GPT-4o baseline.
5.7.1 Comparison with Other Methods
Table 2 compares our Lrq-Fact model with baseline methods. Our proposed approach consistently outperforms all baselines across various datasets, regardless of whether external knowledge is introduced (w/ RAG) or not (w/o RAG). This highlights the effectiveness of incorporating FCQs in enhancing fact-checking performance.
5.7.2 Effect of LLM-Generated FCQs on Fact-Checking Performance
Table 3 reports F1 scores and accuracy (ACC) for various ablation configurations, comparing the baseline GPT-4o model with visual and textual FCQs variations.
Overall Improvements: Incorporating FCQs enhances fact-checking performance across all datasets. For MMFakeBench, the F1 score improves from 49.2 to 71.6 (+45.5%), and accuracy rises from 60.9 to 70.8 (+16.3%). Similarly, for DGM4, the highest F1 score increases from 39.9 to 49.2 (+23.3%), while accuracy improves from 55.9 to 62.3 (+11.5%). Factify also benefits from FCQs integration, with F1 increasing from 72.5 to 75.2 (+3.7%) and accuracy from 71.2 to 73.1 (+2.7%).
Impact of Visual FCQs: Visual FCQs alone yield noticeable improvements in both F1 and accuracy across all datasets. For MMFakeBench, F1 improves by 30.7% (from 49.2 to 64.3), while accuracy rises slightly (+2.0%). A similar trend is observed for DGM4, where F1 improves by 1.3% and accuracy by 2.9%. This improvement can be attributed to the fact that visual FCQs encourage the model to focus more on visual details, prompting a deeper analysis of the image content.
Impact of Textual FCQs without RAG: The addition of textual FCQs without explicit external evidence shows significant gains, especially in DGM4, where the F1 score rises from 39.9 to 47.2 (+18.3%), and accuracy improves by 6.8%. However, this setting is not evaluated for Factify because the task in this dataset is verification rather than detection, and Factify already provides factual reference documents for verification. This improvement can be attributed to the fact that textual FCQs systematically capture all claims within the news and probe their validity. By generating targeted questions, the model breaks down the text into specific factual assertions, allowing for a more structured verification process. This focused approach helps detect inconsistencies, misinterpretations, or misleading statements within the text, ultimately leading to more accurate fact-checking.
Impact of Textual FCQs with RAG: When textual FCQs incorporate external evidence, fact-checking performance further improves. For MMFakeBench, F1 increases to 61.8 (+25.2%), and accuracy improves by 6.2%. In DGM4, this approach achieves the highest F1 (49.4, +23.8%) and accuracy (62.5, +11.8%), highlighting the importance of retrieved evidence. Factify also benefits, with F1 rising to 74.0 (+2.1%) and accuracy to 72.5 (+1.8%).
These results confirm that integrating LLM-generated FCQs, particularly when paired with external evidence, significantly enhances multimodal fact-checking performance.
5.7.3 Effect of Fact-Checking Questions on Fact-Checking Performance
Figure 5: Detailed ablation study result. TFCQs and VFCQs represent Textual FCQs, Visual FCQs respectively.
To further investigate how different types FCQs impact verification performance, we analyze recall scores across different manipulation categories within the MMFakeBench dataset. Figure 5 presents recall values for various settings, including visual FCQs (VFCQs), textual FCQs (TFCQs), and combinations of both with and without RAG.
Impact on Real and Mismatch Cases: While FCQs improve fact-checking performance overall, recall scores for the Real and Mismatch categories decrease. This phenomenon occurs because, in the absence of structured questioning, the model does not apply strict fact-checking criteria and tends to classify most news samples as either Real or Mismatch. However, with FCQs, the model adopts a more expert-like approach, becoming more cautious before confirming a claim as real. This increased scrutiny aligns with human verification processes, reducing false positives but leading to a lower recall in these categories.
Impact on Textual Veracity Distortion (TVD): FCQs significantly enhance recall for cases involving textual manipulation. Without structured questioning, the model relies primarily on its internal knowledge, which may not always be up-to-date or factually accurate. However, incorporating textual FCQs (TFCQs) allows the model to engage in a more structured verification process by utilizing factual information instead of relying solely on pre-trained LLM knowledge. Additionally, integrating external evidence (TFCQs+RAG) enhances the model’s ability to detect inconsistencies in manipulated text, improving misinformation identification. This highlights the value of supplementing LLM-generated responses with retrieved factual data for more reliable fact-checking.
Impact on Visual Veracity Distortion (VVD): Visually manipulated claims pose a challenge for the baseline model, which often lacks the ability to rigorously assess image alterations. However, incorporating visual FCQs (VFCQs) substantially improves recall by prompting the model to analyze visual content more critically. The highest recall is achieved when both VFCQs and TFCQs are used together with external evidence, reinforcing the importance of cross-modal verification.
Across all manipulation types, the most effective setup involves combining VFCQs and TFCQs with external evidence (VFCQs+TFCQs+RAG). While this structured questioning approach makes the model more cautious in verifying real claims, it significantly enhances misinformation detection, ensuring a more reliable fact-checking process. These findings highlight how FCQs improve fact-checking by encouraging a more rigorous verification process. By systematically questioning claims across both textual and visual modalities and incorporating factual retrieval mechanisms, FCQs help the model adopt a more expert-like approach, leading to more precise and reliable fact verification.
5.8 Case Study
To demonstrate the effectiveness of FCQs in detecting and classifying multimodal misinformation, we provide examples in the Appendix (see Figures 7, 8, 9, and 10) showcasing their impact.
6 Conclusion
In this work, we investigate whether LLMs can generate relevant FCQs and whether these LLM-generated FCQs can improve multimodal fact-checking. Through the proposed framework, Lrq-Fact, we demonstrate that LLMs are indeed capable of generating highly relevant and targeted FCQs, effectively addressing a key limitation in AFC systems. Furthermore, our experiments show that incorporating relevant FCQs into the fact-checking process significantly enhances evidence retrieval and improves the overall factuality verification performance. Lrq-Fact outperforms baseline methods, showcasing the effectiveness of FCQ generation in strengthening multimodal fact-checking. These results highlight the potential of LLM-based FCQ formulation as a promising direction in future AFC research, facilitating more reliable and scalable methods.
Limitations
While our study demonstrates that LLM-generated FCQs enhance multimodal fact-checking, several limitations remain. One major limitation is the lack of expert-level validation. Although we evaluate FCQ quality using benchmark datasets and human annotations, we do not assess whether the generated questions exhibit reasoning comparable to domain experts. Incorporating expert evaluations, particularly in specialized fields such as medicine or law, could provide a more rigorous assessment of FCQ quality and alignment with high-quality fact-checking standards.
Another limitation is the absence of a random question baseline. Our experiments compare Lrq-Fact against strong fact-checking baselines, but we do not explicitly test whether randomly generated questions could serve as a control. Introducing a random baseline would help isolate the actual contribution of meaningful FCQ generation from the broader effect of question-driven retrieval. Additionally, our approach is inherently dependent on the capabilities of the underlying LLM. If the model produces vague, misleading, or hallucinated questions, it could negatively impact fact-checking performance by retrieving irrelevant or incorrect evidence. Further investigation into fine-tuned models or more controlled question-generation strategies could mitigate these risks.
Our evaluation process also introduces certain limitations. The FCQs were assessed by PhD students who received detailed instructions and predefined criteria to evaluate question relevance. While this ensures a structured and consistent evaluation process, the annotator pool is relatively small and may not fully represent diverse perspectives. The background knowledge of annotators could influence their judgments, and a broader demographic, including professional fact-checkers or domain experts, may provide a more comprehensive evaluation of FCQ effectiveness.
The reliance on external search engines for evidence retrieval introduces another source of variability. The effectiveness of our approach depends on search engine algorithms, indexing policies, and the availability of high-quality sources, which may not always be consistent. This issue is particularly relevant for fact-checking emerging claims or topics where authoritative sources are scarce. Furthermore, while Lrq-Fact integrates textual and visual evidence for multimodal fact-checking, its reliance on image captions and textual representations of visual content may introduce errors. Exploring direct visual analysis through vision-language models or image embeddings could improve robustness in cases where textual descriptions are insufficient.
Generalization across fact-checking domains remains another open challenge. Although our approach performs well on benchmark datasets, its effectiveness across diverse domains such as political misinformation, scientific fact-checking, and real-time verification is not fully explored. Future work should investigate domain-specific FCQ generation techniques to adapt the framework for specialized fact-checking tasks. Additionally, the computational overhead of our approach may limit its practical deployment. Multiple inference steps, including FCQ generation, evidence retrieval, and multimodal reasoning, contribute to significant resource consumption, making real-time fact-checking more challenging. Optimizing the framework for efficiency would be necessary for large-scale deployment.
Finally, our evaluation process relies on both human annotators and LLM-based scoring to assess FCQ quality. While high agreement between human and model-based evaluations suggests reliability, potential biases in human annotation criteria and systematic artifacts in LLM-generated scoring may still influence results. Future research should explore alternative evaluation metrics and methodologies to ensure robustness and fairness in assessing FCQ effectiveness.
Ethical Statement
This work explores the use of LLMs to generate fact-checking questions (FCQs) for automated verification. While our approach enhances misinformation detection, it raises ethical considerations. LLM-generated FCQs may reflect biases present in training data, potentially influencing fact-checking outcomes. Although we evaluate FCQ quality with human reviewers and benchmark datasets, further efforts are needed to ensure fairness and neutrality.
Another concern is the risk of over-reliance on automation. While LLMs can support fact-checking at scale, they should not replace human judgment, particularly in high-stakes domains like politics and healthcare. Our framework is designed as an assistive tool, reinforcing rather than replacing expert oversight. Privacy considerations are also critical, as our approach retrieves publicly available evidence without processing personally identifiable information. We adhere to ethical data usage practices and fair use policies but recognize the need for continuous alignment with evolving privacy standards.
Finally, transparency and accountability in AI-driven fact-checking remain essential. Black-box decision-making can undermine trust, emphasizing the importance of explainability. By acknowledging these challenges, we advocate for responsible AI deployment, promoting fairness, human oversight, and transparency in automated fact-checking systems.
Appendix A Appendix
A.1 Example Analysis
@@
The questions presented in this case are well-formed and reliable. They are designed to extract key details from both the image and the text, ensuring comprehensive verification. The visual questions effectively ask about the setting and identity, which helps in confirming whether the person and location in the image match the article’s claims. The text-based questions aim to validate the timeline and factual details, ensuring a consistent narrative. These questions are precise and structured to get the best possible answers, making them a solid mechanism for cross-verifying facts.
@@
The questions in this case are reliable and appropriately structured to identify discrepancies between the image and the text. The visual questions ask about the architectural style and contextual clues from the image, while the text-based questions explore the factual accuracy of the claim that this lighthouse is haunted and located in Greece. The questions provide a good framework for fact-checking by encouraging thorough scrutiny of both visual and textual elements. This ensures that any distortions or misrepresentations in the article are effectively highlighted, making the questions a valuable tool for getting to the truth.
@@
The questions are well-crafted to reveal any visual inconsistencies. The visual questions ask about the color and reality of the clock tower, which are key to identifying that the clock tower has been digitally altered. The text-based questions, which probe the existence of such a clock tower in real life, also help uncover discrepancies. These questions are reliable and precise, aimed at extracting the best possible answers and guiding the evaluation of the article’s claims against the evidence provided by the image.
@@
The questions presented are well-designed to highlight the inconsistency between the text and the image. The visual questions ask about the scene and the object the girl is holding, providing clear answers that reveal the mismatch. The text-based questions further confirm this by addressing the article’s lack of accurate description. These questions are well-structured and reliable, allowing for an in-depth examination of both the image and the text to expose cross-modal discrepancies. They guide the analysis toward the best possible answers by focusing on the key elements that need verification.
A.2 Instruct Prompt for Lrq-Fact
@@
Visual Questions Prompt. This stage generates relevant visual questions designed to verify the accuracy, authenticity, and relevance of the visual content in relation to the article. These questions help clarify the image content and assess its relation to the text. The specific prompt for generating these questions is illustrated in Figure 12.
Visual Answers Prompt. After generating the visual questions, this prompt helps in generating answers that analyze the visual content directly from the image. These answers are based on the key elements and actions identified in the image, ensuring that the responses are relevant and insightful. The specific prompt for this is shown in Figure 13.
Textual Questions Prompt. To critically assess the factual claims in the text, this prompt generates relevant questions targeting specific elements such as dates, names, locations, and events. The generated questions aim to challenge the accuracy of the claims made in the article. The specific prompt used for textual questions is shown in Figure 14.
Textual Answers Prompt. After generating textual questions, this prompt enables the model to generate answers using its built-in knowledge. The specific prompt is shown in Figure 15. Additionally, we employ a Retrieval-Augmented Generation (RAG) approach to incorporate factual evidence, ensuring more reliable and verifiable responses. These answers help assess factual accuracy and challenge any unsupported claims in the article. The corresponding RAG-based prompt is illustrated in Figure 16.
Question Quality Assessment Prompt. To evaluate the relevance of generated questions, we use a fact-checking criteria-based prompt that classifies questions as relevant or irrelevant. This assessment considers factors such as alignment with the claim, specificity, and usefulness in verifying factual accuracy. The specific prompt used for this evaluation is illustrated in Figure 17.
@@
A.3 Annotator Details
To evaluate the quality of LLM-generated FCQs, we recruited two PhD students with backgrounds in NLP and computational linguistics. Annotators were provided with detailed instructions and predefined criteria to assess the relevance of each FCQ to the given claim. The evaluation process aimed to ensure consistency and minimize subjectivity in judgment. While this setup provides structured and knowledgeable assessments, the annotator pool is relatively small and may not fully capture diverse perspectives. Future work could incorporate domain experts or professional fact-checkers to further validate FCQ effectiveness across different fact-checking domains.
A.4 Criteria for Evaluating FCQ Quality
To systematically evaluate the quality of LLM-generated fact-checking questions (FCQs), we developed a structured evaluation framework inspired by best practices from established fact-checking methodologies. Our evaluation process consists of two key components: LLM-based assessment and human evaluation, ensuring a rigorous and reliable analysis of question relevance.
Evaluation Framework. We designed our evaluation framework to assess the effectiveness of both visual and textual FCQs. The framework follows ten evaluation criteria, derived from widely accepted fact-checking principles, emphasizing accuracy, credibility, and relevance. These criteria help determine whether the generated FCQs effectively probe factual claims and align with real-world verification standards (Figure 17).
One challenge in evaluating FCQs is ensuring question specificity without over-constraining the verification process. A well-formed FCQ should allow multiple valid answers depending on available evidence while still prompting meaningful fact-checking efforts. Additionally, cross-modal consistency is a key factor in multimodal fact-checking—image-based questions must align with textual claims without introducing unintended biases or assumptions.
LLM-Based vs. Human Evaluation. To ensure consistency, we employ GPT-4o as an automated evaluator, scoring FCQs based on predefined criteria such as logical structure, factual precision, and investigative depth. However, LLM-based evaluations may still miss nuanced contextual ambiguities that a human fact-checker would recognize, such as misleading phrasing or assumptions embedded in a question.
To validate the reliability of the LLM-based assessment, we conducted a human agreement study, comparing GPT-4o’s evaluation results with expert annotations across datasets in Table 1. The goal was to determine the degree of alignment between human and LLM judgments, rather than integrating both assessments into a single process.
Our findings indicate that while LLMs are effective at systematically evaluating FCQs, human reviewers provide valuable qualitative insights, particularly in identifying question formulation errors that could lead to misinformation rather than prevent it.
This subtle difference can impact both retrieval accuracy and the framing of fact-checking results.
Insights from the Evaluation Process.
  • Text-based FCQs generally receive higher relevance scores than image-based FCQs. This discrepancy suggests that LLMs have a better grasp of linguistic verification than visual reasoning, which remains an open challenge in multimodal misinformation detection.
  • Human annotators tend to be stricter in rejecting vague or broad FCQs. LLM-based evaluations show slightly higher acceptance rates for questions that are loosely related to the claim but lack clear fact-checking intent.
  • Context-aware evaluation is critical. Without access to real-world updates, an FCQ might appear factually valid but be outdated or misleading in light of new developments. This highlights the importance of external knowledge retrieval in automated fact-checking pipelines.
By comparing human and LLM-based assessments, our study confirms that GPT-4o produces highly relevant FCQs with near-human accuracy. However, human reviewers remain essential in refining question design and identifying subtle logical inconsistencies that automated evaluations may overlook.
Criteria Definition
Critical Thinking and Skepticism Challenges assumptions, probes deeper into claims, avoids taking information at face value.
Analytical Depth Breaks down complex statements into verifiable components.
Systematic Approach Follows a structured methodology in assessing sources, claims, and evidence.
Precision & Specificity Clear, direct, and free from vague or overly broad wording.
Factual Accuracy Focuses on verifying evidence, checking primary sources, and detecting misinformation.
Logical Consistency Identifies contradictions, misleading narratives, or inconsistencies.
Source Credibility & Bias Detection Evaluates the reliability of cited sources and potential biases.
Context Awareness Considers the broader context surrounding the claim.
Comparative Thinking Encourages cross-referencing with established facts or alternative perspectives.
Repeatability and Objectivity Can be applied consistently across different cases without personal bias.
Table 4: Criteria for assessing the accuracy, credibility, and reliability of FCQs.
A.5 Dataset Descriptions and Details
To assess the effectiveness of LLM-generated fact-checking questions in multimodal misinformation detection, we utilize three benchmark datasets: MMFakeBench, DGM4, and Factify. These datasets encompass a wide range of real and manipulated image-text pairs, enabling a comprehensive evaluation of textual, visual, and cross-modal inconsistencies. Each dataset provides a distinct annotation scheme, capturing various types of misinformation, from textual distortions to manipulated images and multimodal inconsistencies.
MMFakeBench Dataset. This dataset serves as a benchmark for multimodal misinformation detection. It categorizes misinformation into three primary types:
  • Textual Veracity Distortion (TVD): Fake or misleading textual claims.
  • Visual Veracity Distortion (VVD): Manipulated or AI-generated images.
  • Cross-Modal Consistency Distortion (CMM): Mismatches between text and images.
Each sample in MMFakeBench is annotated based on:
  • Whether the claim text is factually correct.
  • Whether the accompanying image has been manipulated.
  • Whether the text-image pair is consistent or inconsistent.
The dataset provides a structured framework to evaluate misinformation detection across multiple manipulation types, incorporating diverse real-world scenarios.
DGM4 Dataset. The DGM4 (Grounding Multi-Modal Media Manipulation) dataset is a large-scale collection of manipulated and real news samples focusing on human-centric content. It contains approximately 230,000 samples, distributed as follows:
  • 77,426 pristine (real) image-text pairs.
  • 152,574 manipulated samples, generated using:
    • Face Swap (FS): 66,722 samples.
    • Face Attribute Manipulation (FA): 56,411 samples.
    • Text Swap (TS): 43,546 samples.
    • Text Attribute Manipulation (TA): 18,588 samples.
    • Mixed Manipulation Pairs: 32,693 samples combining text and image edits.
Factify Dataset. This dataset is a multimodal fact verification dataset containing 50,000 samples collected from Twitter and online news sources in the United States and India. Each sample consists of:
  • Claim text: A short statement, often extracted from tweets.
  • Claim image: The corresponding image that either supports or contradicts the claim.
  • OCR text: Extracted text from the claim image.
  • Document text: A news article serving as supporting evidence.
  • Document image: An image from the referenced news article.
Samples in Factify are categorized into five classes:
  • Support_Text: The claim text is supported by the document text, but images are dissimilar.
  • Support_Multimodal: Both the claim text and image match the document text and image.
  • Insufficient_Text: The document does not provide enough textual evidence to support or refute the claim.
  • Insufficient_Multimodal: The document image matches the claim image, but the text lacks confirmation.
  • Refute: The document contradicts both the claim text and the claim image.
The dataset provides a benchmark for multimodal fact verification, leveraging both textual and visual evidence to assess claim veracity.
Figure 6: The overview pipeline of our Lrq-Fact framework consists of four key components: (a) Image Description, which provides detailed contextual descriptions of the image; (b) Visual FCQs, aimed at assessing the accuracy of the visual content; and (c) Textual FCQs, which detect textual inaccuracies, contradictions, or unsupported claims. Finally, all the gathered information is synthesized in (d) the Final Judgment & Explanation module, where a rule-based decision-maker generates both the prediction results and comprehensive explanations.